OpenStudy (anonymous):

divide 7+9i/8+9i. It has to be simplified

OpenStudy (anonymous):

7 + ((9 * i) / 8) + (9 * i) = 7 + 10.12500 i

OpenStudy (anonymous):

\(\huge \color{blue}{ \frac{a+b.i}{c+d.i}=\frac{(a+b.i)(c-d.i)}{(c+d.i)(c-d.i)}=\frac{(a+b.i)(c-d.i)}{c^2+b^2}}\)